OpenStudy (anonymous):

Divide. Write the answer with positive exponents only. 32a^2b^2/8ab^4

OpenStudy (anonymous):

\[\color{red}{\frac{a^x}{a^y} = a^{x-y}}\]

OpenStudy (anonymous):

For example : \[\frac{a^5}{a^2} = a^{5-2} = a^3\]

OpenStudy (anonymous):

Now: the same is to do with your question: \[\frac{32\cdot a^2 \cdot b^2}{8 \cdot a \cdot b^4} = ??\]

OpenStudy (anonymous):

Solve one by one : \[\frac{32}{8} = ??\]

OpenStudy (anonymous):

Now a : \[\frac{a^2}{a} = ??\]

OpenStudy (anonymous):

Now for b : \[\frac{b^2}{b^4} = ??\]

OpenStudy (anonymous):

Are you getting ??

OpenStudy (anonymous):

Are you interested in replying @maryam123 ??